OpenStudy (anonymous):

A solution is made by dissolving 4.0 moles of sodium chloride (NaCl) in 2.05 kilograms of water. If the molal boiling point constant for water (Kb) is 0.51 °C/m, what would be the boiling point of this solution? Show all of the work needed to solve this problem.

OpenStudy (anonymous):

\[\Delta T = Kb * b\] DT= 0,51 * (4 mol / 2,05 kg) = 0,995 °C so boiling point of solution would be 100+0,995 = 100,995 °C